# MaxK-GNN: Towards Theoretical Speed Limits for Accelerating Graph Neural Networks Training

### 22 Feb 2024 v3

@@ Abstract.
We conduct extensive evaluations of MaxK-GNN and report the end-to-end system run-time. Experiments show that MaxK-GNN system could approach the speedup limit according to Amdahl’s law. We achieve comparable accuracy to SOTA GNNs, but at a significantly increased speed: 3.22$\times$/4.24$\times$ speedup (vs. 5.52$\times$/7.27$\times$) on Reddit compared to DGL and GNNAdvisor implementations. Our implementation can be found on GitHub [^1] https://github.com/harveyp123/MaxK-GNN.
@@ 1. Introduction
We conduct extensive evaluations of MaxK-GNN system within the context of a single-GPU, full-batch, GNN training workload. We report the end-to-end system run-time rather than floating point operations per second (FLOPS) analysis. Experiments show that our MaxK-GNN system could approach the speedup limit in kernel design according to Amdahl’s law (gustafson1988reevaluating, ). The performance gaps between our results and the limits, e.g., 3.22$\times$/4.24$\times$ compared to 5.52$\times$/7.27$\times$ for Reddit dataset using MaxK-GNN with GraphSAGE, is from the accumulation stage of SpGEMM and dense row prefetching stage of SSpMM, which empirically are difficult to further optimize, currently.
@@ 5. Evaluation
We offer a comprehensive assessment of MaxK-GNN, highlighting its performance advantages and applicability in the broader context of graph-based learning and computation. Our evaluation strategy begins with an in-depth analysis of both the forward SpGEMM and the backward SSpMM kernels. We carefully assess these components under a range of $K$ values, comparing the performance with existing implementations such as SpMM, as found in GNNAdvisor (wang2021gnnadvisor, ) and cuSPARSE v12.0 (naumov2010cusparse, ). Following this detailed kernel-level examination, we present our end-to-end evaluation. This encapsulates both the accuracy and speedup metrics to showcase the efficacy of the MaxK-GNN framework in addressing general graph learning problems.
Figure 9. End-to-end MaxK-GNN system evaluation for GraphSAGE (hamilton2017inductive, ), GCN (kipf2016semi, ), and GIN (xu2018gin, ) models and Reddit, ogbn-proteins, ogbn-produces, Yelp, and Flickr datasets. MaxK nonlinearity setting: $k=[2,4,8,16,32,64,96,128,192]$. We compare the end to end training speedup over DGL with cuSPARSE (dgl_amazon, ) framework and GNNAdvisor (wang2021gnnadvisor, ). spd: speedup.
@@ 5.1. Experimental Setup
For end-to-end MaxK-GNN system evaluation, we benchmark five datasets that range from small to medium-scale graphs. Specifically, the selected datasets are Flickr (JulianFlickr, ), for the categorization of image types based on descriptors and shared attributes; Yelp (Zengiclr20, ), for the classification of user-generated reviews pertaining to businesses and services; Reddit (reddit, ), for community prediction using posts’ content and users’ comments; ogbn-products (hu2020open, ), for Amazon product classification via customer reviews; and ogbn-proteins (hu2020open, ), for the prediction of protein function presence. This suite of datasets not only covers a broad spectrum of applications but also facilitates the understanding of how MaxK-GNN performs under various scenarios.
@@
For the end-to-end evaluation of the MaxK-GNN system, we integrate MaxK nonlinearity with three graph models: GraphSAGE (hamilton2017inductive, ), GCN (kipf2016semi, ), and GIN (xu2018gin, ). The GIN model (xu2018gin, ) is noteworthy for its unique aggregation function, serving as a reference for advanced GNNs such as Graph Attention Networks (GAT) (GATConv, ). To ensure fairness, all models are trained in full batch graph learning mode with the MEAN aggregator for GraphSAGE. The detailed parameter settings of MaxK-GNN are given in Table 3. Our evaluation spans $k=[2,4,8,16,32,64,96,128,192]$ to find the best trade-off between model accuracy and system speedup, highlighting the versatility of our approach.
Environment Setup. We construct the MaxK-GNN by implementing the SpGEMM & SSpMM kernels using C++ and CUDA C. The front-end is built with Python/Pytorch to provide a user-friendly interface. Our principal evaluation platform consists of a high-performance server equipped with 32-core AMD EPYC 7513 CPU and NVIDIA A100 80GB GPU (choquette2021nvidia, ). This computing environment is utilized for both kernel-level and system-level evaluations.
@@ 5.2. MaxK-GNN Kernel Evaluation
The backward SSpMM kernel fully exploits the global memory access coalescing and traffic reduction by combining the outer product-based approach with dense row prefetching. The SSpMM kernel achieves better speedup performance than the forward SpGEMM kernel. For graphs with average degrees greater than 50, the average speedup of the SSpMM kernel at $k=8,16,32,64$ is 6.93×6.93$\times$6.93 ×, 5.39×5.39$\times$5.39 ×, 2.55×2.$5$, 1.46×1.46$\times$1.46 × respectively, as compared to the cuSPARSE (naumov2010cusparse, ) and 9.57×9.57$\times$9.57 ×, 7.46×7.46$\times$7.46 ×, 3.55×3.$5$, 2.04×2.04$\times$2.04 ×, respectively, as compared to the GNNAdvisor (wang2021gnnadvisor, ). Given that a MaxK-GNN layer training pipeline requires computing both forward SpGEMM and the backward SSpMM, the SSpMM kernel speedup can also benefit the end-to-end training speedup. When $k\leq 128$, the SSpMM kernel can effectively bring speedup to 87.5% of all test cases compared to cuSPARSE (naumov2010cusparse, ), and 98.4% of all test cases compared to GNNAdvisor (wang2021gnnadvisor, ). It also exhibits significant versatility.
5.3. End-to-end MaxK-GNN Evaluation
@@ 6. Conclusion and Future Work
Recent advancements in MLaaS (Machine Learning as a Service) have primarily centered on accelerating inference on various platforms (xie2023accel, ; peng2023autorep, ; chen2023recontab, ; medusa, ; thorat2023advanced, ; peng2022towards, ; chen2019claims, ; peng2023lingcn, ; chen2018data, ; peng2023rrnet, ; wu2022optimizing, ; peng2022length, ; yan2023improving, ; peng2023pasnet, ; sheng2023muffin, ). In this paper, we present MaxK-GNN, a high-performance GPU training system integrating algorithm and system innovation. (i) We introduce the MaxK nonlinearity and provide a theoretical analysis of MaxK nonlinearity as a universal approximator, and present the Compressed Balanced Sparse Row (CBSR) format, designed to store the data and index of the feature matrix after nonlinearity; (ii) We design a coalescing enhanced forward computation with row-wise product-based SpGEMM Kernel using CBSR for input feature matrix fetching and strategic placement of a sparse output accumulation buffer in shared memory; (iii) We develop an optimized backward computation with outer product-based and SSpMM Kernel. Experiments show that our MaxK-GNN system could approach the theoretical speedup limit according to Amdahl’s law. We achieve comparable accuracy to existing GNNs, but at a significantly increased speed: 3.22$\times$/4.24$\times$ speedup (vs. theoretical limits, 5.52$\times$/7.27$\times$) on Reddit.
@@ 7. Acknowledgments
This research was supported in part by the Northeastern University Institute for Experiential AI, the NSF IUCRC Center for Hardware and Embedded Systems Security and Trust (CHEST), NSF CNS-2247893, the Semiconductor Research Corporation (SRC) Artificial Intelligence Hardware program, and Advanced Micro Devices (AMD).

### 26 Feb 2024 v4

# MaxK-GNN: Extremely Fast GPU Kernel Design for Accelerating Graph Neural Networks Training

@@ Abstract.
We conduct extensive evaluations of MaxK-GNN and report the system training time. Experiments show that MaxK-GNN system could approach the speedup limit according to Amdahl’s law. We achieve comparable accuracy to SOTA GNNs, but at a significantly increased speed: 3.22$\times$/4.24$\times$ speedup (vs. 5.52$\times$/7.27$\times$) on Reddit compared to DGL and GNNAdvisor implementations. Our implementation can be found on GitHub [^1] https://github.com/harveyp123/MaxK-GNN.
@@
journalyear: 2024copyright: acmlicensedconference: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; April 27-May 1, 2024; La Jolla, CA, USAbooktitle: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’24), April 27-May 1, 2024, La Jolla, CA, USAdoi: 10.1145/3620665.3640426isbn: 979-8-4007-0385-0/24/04
@@ 1. Introduction
We conduct extensive evaluations of MaxK-GNN system within the context of a single-GPU, full-batch, GNN training workload. We report the system training time rather than floating point operations per second (FLOPS) analysis. Experiments show that our MaxK-GNN system could approach the speedup limit in kernel design according to Amdahl’s law (gustafson1988reevaluating, ). The performance gaps between our results and the limits, e.g., 3.22$\times$/4.24$\times$ compared to 5.52$\times$/7.27$\times$ for Reddit dataset using MaxK-GNN with GraphSAGE, is from the accumulation stage of SpGEMM and dense row prefetching stage of SSpMM, which empirically are difficult to further optimize, currently.
@@ 5. Evaluation
We offer a comprehensive assessment of MaxK-GNN, highlighting its performance advantages and applicability in the broader context of graph-based learning and computation. Our evaluation strategy begins with an in-depth analysis of both the forward SpGEMM and the backward SSpMM kernels. We carefully assess these components under a range of $K$ values, comparing the performance with existing implementations such as SpMM, as found in GNNAdvisor (wang2021gnnadvisor, ) and cuSPARSE v12.0 (naumov2010cusparse, ). Following this detailed kernel-level examination, we present our system training time evaluation. This encapsulates both the accuracy and speedup metrics to showcase the efficacy of the MaxK-GNN framework in addressing general graph learning problems.
Figure 9. MaxK-GNN system evaluation for GraphSAGE (hamilton2017inductive, ), GCN (kipf2016semi, ), and GIN (xu2018gin, ) models and Reddit, ogbn-proteins, ogbn-produces, Yelp, and Flickr datasets. MaxK nonlinearity setting: $k=[2,4,8,16,32,64,96,128,192]$. We compare the end to end training speedup over DGL with cuSPARSE (dgl_amazon, ) framework and GNNAdvisor (wang2021gnnadvisor, ). spd: speedup.
@@ 5.1. Experimental Setup
For MaxK-GNN system evaluation, we benchmark five datasets that range from small to medium-scale graphs. Specifically, the selected datasets are Flickr (JulianFlickr, ), for the categorization of image types based on descriptors and shared attributes; Yelp (Zengiclr20, ), for the classification of user-generated reviews pertaining to businesses and services; Reddit (reddit, ), for community prediction using posts’ content and users’ comments; ogbn-products (hu2020open, ), for Amazon product classification via customer reviews; and ogbn-proteins (hu2020open, ), for the prediction of protein function presence. This suite of datasets not only covers a broad spectrum of applications but also facilitates the understanding of how MaxK-GNN performs under various scenarios.
@@
For the system training time evaluation of MaxK-GNN, we integrate MaxK nonlinearity with three graph models: GraphSAGE (hamilton2017inductive, ), GCN (kipf2016semi, ), and GIN (xu2018gin, ). The GIN model (xu2018gin, ) is noteworthy for its unique aggregation function, serving as a reference for advanced GNNs such as Graph Attention Networks (GAT) (GATConv, ). To ensure fairness, all models are trained in full batch graph learning mode with the MEAN aggregator for GraphSAGE. The detailed parameter settings of MaxK-GNN are given in Table 3. Our evaluation spans $k=[2,4,8,16,32,64,96,128,192]$ to find the best trade-off between model accuracy and system speedup, highlighting the versatility of our approach.
Environment Setup. We construct the MaxK-GNN by implementing the SpGEMM & SSpMM kernels using C++ and CUDA C. Our principal evaluation platform consists of a high-performance server equipped with 32-core AMD EPYC 7513 CPU and NVIDIA A100 80GB GPU (choquette2021nvidia, ). This computing environment is utilized for both kernel-level and system-level evaluations.
@@ 5.2. MaxK-GNN Kernel Evaluation
The backward SSpMM kernel fully exploits the global memory access coalescing and traffic reduction by combining the outer product-based approach with dense row prefetching. The SSpMM kernel achieves better speedup performance than the forward SpGEMM kernel. For graphs with average degrees greater than 50, the average speedup of the SSpMM kernel at $k=8,16,32,64$ is 6.93×6.93$\times$6.93 ×, 5.39×5.39$\times$5.39 ×, 2.55×2.$5$, 1.46×1.46$\times$1.46 × respectively, as compared to the cuSPARSE (naumov2010cusparse, ) and 9.57×9.57$\times$9.57 ×, 7.46×7.46$\times$7.46 ×, 3.55×3.$5$, 2.04×2.04$\times$2.04 ×, respectively, as compared to the GNNAdvisor (wang2021gnnadvisor, ). Given that a MaxK-GNN layer training pipeline requires computing both forward SpGEMM and the backward SSpMM, the SSpMM kernel speedup can also benefit the training speedup. When $k\leq 128$, the SSpMM kernel can effectively bring speedup to 87.5% of all test cases compared to cuSPARSE (naumov2010cusparse, ), and 98.4% of all test cases compared to GNNAdvisor (wang2021gnnadvisor, ). It also exhibits significant versatility.
5.3. MaxK-GNN System Training Time Evaluation
@@ 6. Conclusion and Future Work
Recent advancements in MLaaS (Machine Learning as a Service) have primarily centered on accelerating inference on various platforms (xie2023accel, ; peng2023autorep, ; chen2023recontab, ; medusa, ; thorat2023advanced, ; peng2022towards, ; chen2019claims, ; peng2023lingcn, ; chen2018data, ; peng2023rrnet, ; wu2022optimizing, ; peng2022length, ; yan2023improving, ; peng2023pasnet, ; sheng2023muffin, ). In this paper, we present MaxK-GNN, a high-performance GPU training system integrating algorithm and system innovation. (i) We introduce the MaxK nonlinearity and provide a theoretical analysis of MaxK nonlinearity as a universal approximator, and present the Compressed Balanced Sparse Row (CBSR) format, designed to store the data and index of the feature matrix after nonlinearity; (ii) We design a coalescing enhanced forward computation with row-wise product-based SpGEMM Kernel using CBSR for input feature matrix fetching and strategic placement of a sparse output accumulation buffer in shared memory; (iii) We develop an optimized backward computation with outer product-based and SSpMM Kernel. Experiments show that our MaxK-GNN system could approach the limit according to Amdahl’s law. We achieve comparable accuracy to existing GNNs, but at a significantly increased speed: 3.22$\times$/4.24$\times$ speedup (vs. theoretical limits, 5.52$\times$/7.27$\times$) on Reddit.
@@ 7. Acknowledgments
This research was supported in part by the Northeastern University Institute for Experiential AI, the NSF IUCRC Center for Hardware and Embedded Systems Security and Trust (CHEST), NSF SHF-2340273, the Semiconductor Research Corporation (SRC) Artificial Intelligence Hardware program, and Advanced Micro Devices (AMD).